# Harmonic Self-Conditioned Flow Matching for joint Multi-Ligand Docking and Binding Site Design

### v3

@@ 2 Related Work
Other applications of flow matching to biomolecular problems include generating Boltzmann distributions of small molecules Klein et al. (2023), protein structure generation (Yim et al., 2023; Bose et al., 2023) and small molecule generation (Song et al., 2023). We explain flow matching in Section 3.1.
@@ 3.1 HarmonicFlow Structure Generation
Harmonic Prior. Any prior can be used for $p_{0}$ in the flow matching framework. We choose a harmonic prior as in EigenFold (Jing et al., 2023) that samples atoms to be close to each other if they are connected by a bond. We identify this as an especially valuable inductive bias when dealing with multiple molecules and ions since atoms of different molecules are already spatially separated at $t=0$ as visualized in Figure 3.
This prior is constructed based on covalent bonds that define a graph with adjacency matrix ${\bm{A}}$ from which we can construct the graph Laplacian ${\bm{L}}={\bm{D}}-{\bm{A}}$ where ${\bm{D}}$ is the degree matrix. The harmonic prior is then $p_{0}({\bm{x}}_{0})\propto exp(-\frac{1}{2}{\bm{x}}_{0}^{T}{\bm{L}}{\bm{x}}_{0})$ which can be sampled as a transformed gaussian.
Structure Self-conditioning. With this, we aim to bring AlphaFold2’s (Jumper et al., 2021) successful recycling strategy to flow models for structure generation. Recycling enables training a deeper structure predictor without additional memory cost by performing multiple forward passes while only computing gradients for the last. For flow matching, we achieve the same by adapting the discrete diffusion model self-conditioning approach of Chen et al. (2023) as follows:
@@ 3.2 FlowSite Binding Site Design
Fake Ligand Data Augmentation. This strategy is based on the evidence of Polizzi & DeGrado (2020) that a protein’s sidechain-sidechain interactions are similar to sidechain-ligand interactions for tight binding. In our optional data augmentation, we train with 20% of the samples having a ”fake ligand”. Given a protein, we construct a fake ligand as the atoms of a randomly selected residue that has at least 4 other residues within 4Å heavy atom distance. Additionally, we modify the protein by removing the residue that was chosen as the fake ligand and the residues that are within 7 positions next to that residue in the protein chain, as visualized in Figure 5.
@@ 3.3 Architecture
Equivariant refinement Tensor Field Network (TFN) layers. Based on TFN (Geiger et al., 2020), these layers capture the important inductive bias of SE(3)-equivariance (translating and rotating the input will equally translate and rotate the output). They are a remarkably simple yet effective tweak from previous message passing TFNs (Jing et al., 2022; Corso et al., 2023), where we instead update and refine ligand coordinates with each layer akin to EGNNs (Hoogeboom et al., 2022).
@@ 4 Experiments
We evaluate FlowSite with multiple datasets, splits, and settings. Every reported number is averaged over 10 generated samples for each ligand. Precise experimental details are in Appendix E and code to reproduce each experiment is at https://anonymous.4open.science/r/wolf. The main questions we seek to answer with the experiments are:
@@ 4.2 Question 1: HarmonicFlow Structure Generation
PDBBind docking results. In Table 1, we find that our flow matching based HarmonicFlow outperforms product-space diffusion in average sample quality. The sampled conformations in Figure 12 and videos of the generation process show that HarmonicFlow produces chemically plausible structures and well captures the physical constraints of interatomic interactions without the need to restrict conformational flexibility to torsion angles. No separate losses, rotation updates, or expensive torsion angle updates are required - HarmonicFlow is arguably a cleaner and simpler solution. Thus, it is a promising future direction to further explore HarmonicFlow for docking and other biomolecular structure generation tasks.
@@ 4.3 Question 2: FlowSite Binding Site Recovery
Metrics. We use two metrics, sequence recovery and our new BLOSUM score. Sequence recovery is the percentage of generated residue types of the contact residues (those with a heavy atom within 4Å of the ligand) that are the same as in the original binding site. For designing new sequences via inverse folding, this metric has been shown to correlate well with more expensive computational metrics and experimental performance (Dauparas et al., 2022). However, this metric only rewards exact matches, it cannot account for multiple correct solutions, and there is no notion of amino acid similarity encoded in it. Having lower penalties for mismatches of very similar amino acids would be more meaningful. To address this, we propose BLOSUM score, which takes evolutionary and physicochemical similarity into account. A precise definition is in Appendix A.1.
Metrics that correlate with binding affinity, such as MM-PBSA or docking scores of traditional docking software, usually require the atomic structure of the side chains as input, which is not available. Furthermore, they were only validated for discriminating different ligands, while in our design task, we desire to discriminate different binding sites.
Baselines. PiFold (no ligand) is the architecture of Gao et al. (2023a) and does not use any ligand information. In PiFold (2D ligand), we first process the ligand with PNA (Corso et al., 2020) message passing and pass its features as additional input to the PiFold architecture. Ground Truth Pos and Random Ligand Pos use the architecture of FlowSite without the ligand structure prediction layers. Instead, the ligand positions are either the ground truth bound structure or sampled from a standard Normal at the pocket’s alpha carbon center of mass. Similarly, DiffDock-Pocket Pos uses fixed positions and the same architecture, but the positions are given by the pocket-level docking tool DiffDock-Pocket (Plainer et al., 2023) (we only have results for single ligands since it is not implemented for multi-ligands). The oracle Ground Truth Pos method also uses fake ligand data augmentation.
@@ 5 Conclusion
6 Impact Statement
@@ Appendix E Experimental Setup Details
In this section, we provide additional details on how our experiments were run next to the exact commands and code to reproduce the results available at https://anonymous.4open.science/r/wolf. In all of the paper, we only consider heavy atoms (no hydrogens).

### v4

@@ 2 Related Work
Other applications of flow matching to biomolecular problems include generating Boltzmann distributions of small molecules Klein et al. (2023), protein structure generation (Yim et al., 2023a; Bose et al., 2023) and small molecule generation (Song et al., 2023). We explain flow matching in Section 3.1.
@@ 3.1 HarmonicFlow Structure Generation
Harmonic Prior. Any prior can be used for $p_{0}$ in the flow matching framework. We choose a harmonic prior as in EigenFold (Jing et al., 2023) that samples atoms to be close to each other if they are connected by a bond. Potentially, this inductive bias is especially helpful when dealing with multiple molecules and ions since atoms of different molecules are already spatially separated at $t=0$ as visualized in Figure 3.
This prior is constructed based on the ligand’s covalent bonds that define a graph with adjacency matrix ${\bm{A}}$ from which we can construct the graph Laplacian ${\bm{L}}={\bm{D}}-{\bm{A}}$ where ${\bm{D}}$ is the degree matrix. The harmonic prior is then $p_{0}({\bm{x}}_{0})\propto exp(-\frac{1}{2}{\bm{x}}_{0}^{T}{\bm{L}}{\bm{x}}_{0})$ which can be sampled as a transformed gaussian.
Structure Self-conditioning. With this, we aim to bring AlphaFold2’s (Jumper et al., 2021) successful recycling strategy to flow models for structure generation. Recycling enables training a deeper structure predictor without additional memory cost by performing multiple forward passes while only computing gradients for the last. For flow matching, we achieve the same by adapting the discrete diffusion model self-conditioning approach of Chen et al. (2023), similar to self-conditioning in protein structure generation diffusion models (Yim et al., 2023b; Watson et al., 2023).
@@ 3.2 FlowSite Binding Site Design
Fake Ligand Data Augmentation. This strategy is based on the evidence of Polizzi & DeGrado (2020) that a protein’s sidechain-sidechain interactions are similar to sidechain-ligand interactions for tight binding. In our optional data augmentation, we train with 20% of the samples having a ”fake ligand”. Given a protein, we construct a fake ligand as the atoms of a randomly selected residue that has at least 4 other residues within 4Å heavy atom distance. Additionally, we modify the protein by removing the residue that was chosen as the fake ligand and the residues that are within 7 positions next to that residue in the protein chain (Figure 5). This data augmentation was also employed in concurrent work (Dauparas et al., 2023; Corso et al., 2024).
@@ 3.3 Architecture
Equivariant refinement TFN layers. Based on Tensor Field Networks (TFN) (Geiger et al., 2020), these layers are a simple yet effective tweak from previous use cases of message passing TFNs (Jing et al., 2022; Corso et al., 2023), where we instead update and refine ligand coordinates with each layer akin to EGNNs (Hoogeboom et al., 2022). See Appendix A.5 for more details.
@@ 4 Experiments
We evaluate FlowSite with multiple datasets, splits, and settings. Every reported number is averaged over 10 generated samples for each ligand. Precise experimental details are in Appendix E and code to reproduce each experiment is at https://github.com/HannesStark/FlowSite. The main questions we seek to answer with the experiments are:
@@ 4.2 Question 1: HarmonicFlow Structure Generation
PDBBind docking results. In Table 1, we find that our flow matching based HarmonicFlow outperforms product-space diffusion in average sample quality. The sampled conformations in Figure 12 show that HarmonicFlow produces chemically plausible structures and well captures the physical constraints of interatomic interactions without the need to restrict conformational flexibility to torsion angles. No separate losses, rotation updates, or expensive torsion angle updates are required - HarmonicFlow is arguably a cleaner and simpler solution. Thus, it is a promising future direction to further explore HarmonicFlow for docking and other biomolecular structure generation tasks.
@@ 4.3 Question 2: FlowSite Binding Site Recovery
Metrics. We use two metrics, sequence recovery and our BLOSUM score. Sequence recovery is the percentage of generated residue types of the contact residues (those with a heavy atom within 4Å of the ligand) that are the same as in the original binding site. This metric only rewards exact matches, it cannot account for multiple correct solutions, and there is no notion of amino acid similarity encoded in it. Having lower penalties for mismatches of very similar amino acids would be more meaningful. To address this, we propose BLOSUM score, which takes evolutionary and physicochemical similarity into account. A precise definition is in Appendix A.1.
We note that metrics that correlate with binding affinity, such as MM-PBSA or docking scores of traditional docking software, usually require the atomic structure of the side chains as input, which is not available. Furthermore, they were only validated for discriminating different ligands, while in our design task, we desire to discriminate different binding sites.
Baselines. PiFold (no ligand) is the architecture of Gao et al. (2023a) and does not use any ligand information and ProteinMPNN (no ligand) is the analog with the architecture of (Dauparas et al., 2022). In PiFold (2D ligand), we first process the ligand with PNA (Corso et al., 2020) message passing and pass its features as additional input to the PiFold architecture. Ground Truth Pos and Random Ligand Pos use the architecture of FlowSite without the ligand structure prediction layers. Instead, the ligand positions are either the ground truth bound structure or sampled from a standard Normal at the pocket’s alpha carbon center of mass. Similarly, DiffDock-Pocket Pos uses fixed positions and the same architecture, but the positions are given by the pocket-level docking tool DiffDock-Pocket (Plainer et al., 2023) (we only have results for single ligands since it is not implemented for multi-ligands). The oracle Ground Truth Pos method also uses fake ligand data augmentation.
@@ 5 Conclusion
Acknowledgments
We thank Jason Yim, Gabriele Corso, Rachel Wu, Felix Faltings, Jeremy Wohlwend, MinGyu Choi, Juno Nam, Soojung Yang, Nick Polizzi, Jody Mou, Michael Plainer, Julia Balla, Shangyuan Tong, and Peter Holderrieth for insightful discussions and feedback.
This work was supported by the NSF Expeditions grant (award 1918839: Collaborative Research: Understanding the World Through Code), the Machine Learning for Pharmaceutical Discovery and Synthesis (MLPDS) consortium, the Abdul Latif Jameel Clinic for Machine Learning in Health, the DTRA Discovery of Medical Countermeasures Against New and Emerging (DOMANE) threats program, the DARPA Accelerated Molecular Discovery program, the NSF AI Institute CCF-2112665, the NSF Award 2134795, and the GIST-MIT Research Collaboration grant.
Impact Statement
@@ Appendix E Experimental Setup Details
In this section, we provide additional details on how our experiments were run next to the exact commands and code to reproduce the results available at https://github.com/HannesStark/FlowSite. In all of the paper, we only consider heavy atoms (no hydrogens).